An object travels around a circle 10.0 full turns in 2.5 seconds. Calculate the angular displacement , theta in radians


1
Expert's answer
2022-01-17T09:54:48-0500
"\rm angular\; displacement=number\; of\; turns\times 2\pi"

"\theta=2\pi N=2\pi *10=20\pi \:\rm rad=62.8\: rad"